OpenStudy (anonymous):

A 8.048 gram sample of an organic compound containing C, H and O is analyzed by combustion analysis and 12.00 grams of CO2 and 3.684 grams of H2O are produced. In a separate experiment, the molar mass is found to be 118.1 g/mol. Determine the empirical formula and the molecular formula of the organic compound.

OpenStudy (anonymous):

Atomic weights: H: 1 C: 12 O: 16 I'm calling the mystery compound CHO? Combustion analysis adds O2, which is where the extra mass on the consumption side comes from. 8.048g CHO? + 7.636g O2 -> 12.00g CO2 + 3.684g H2O (12.00 + 3.684 - 8.048 = 7.636) 1mol H2O is 18g: 2g H + 16g O 3.684g H2O is produced 3.684 * (2/18) = 0.4122 g H 3.684 * (16/18) = 3.272g O 1mol CO2 is 44g: 12g C + 32g O 12.00g CO2 is produced 12.00 * (12/44) = 3.275g C 12.00 * (32/44) = 8.725g O 1mol O2 is 32g: 32g O 7.636g O2 is consumed (all from O) The O provided by CHO? is 3.272 + 8.725 - 7.636 = 4.361g. All the C and H produced only came from the CHO? so that's easy to determine. 8.048g CHO? contains: 0.4122g H 3.275g C 4.361g O But what portion of the 118.1g/mol do each of those represent? (0.4122/8.048) * 118.1 = 6.05g H/mol CHO? = 6 H atoms (1g/atom) (3.275/8.048) * 118.1 = 48g C = 4 C atoms (12g/atom) (4.361/8.048) * 118.1 = 64g O = 4 O atoms (16g/atom) So the molecular formula is C4H6O4. The empirical formula is C2H3O2. According to http://www.nauticus.org/chemistry/chememperical.html.